OpenStudy (sbentleyaz):

For which pair of functions f(x) and g(x) below will the lim x-> infinity f(x)g(x) NOT EQUAL 0 f(x) = 10x + e-x; g(x) = f(x) = x2; g(x) = e-4x f(x) =(Lnx)3; g(x) = f(x) = ; g(x) = e-x

OpenStudy (sbentleyaz):

I think that a or c could work so how do you determine which (if thats even correct)

OpenStudy (sbentleyaz):

I graphed them all

OpenStudy (solomonzelman):

Do you know the L'Hospital's rule?

OpenStudy (sbentleyaz):

and a and c are the only ones that do not approach 0

OpenStudy (sbentleyaz):

not really

OpenStudy (solomonzelman):

For a, we would set \(\color{black}{\displaystyle \lim_{x\to\infty}(10x)(e^{-x})=\lim_{x\to\infty}\frac{10x}{e^x} }\)

OpenStudy (solomonzelman):

You don't even need the L'Hospital's rule, you can just note that from roughly \(x=6\) and on, \(2^x>10x\), so a larger sequence would be \(2^x/e^x=(2/e)^x\).

OpenStudy (solomonzelman):

Well, we know that \(\color{black}{\displaystyle \lim_{x\to\infty}\frac{2^x}{e^x}=0 }\), and therefore, since \(\color{black}{\displaystyle \frac{10x}{e^x}<\frac{2^x}{e^x} }\), we can note that \(\color{black}{\displaystyle \lim_{x\to\infty}\frac{10x}{e^x}=0 }\). (This is without L'H'S, although with it, it would be a matter of seconds to find the limit.)

OpenStudy (solomonzelman):

What is g(x) in c?

OpenStudy (sbentleyaz):

why did you divide for 10x/e^x

OpenStudy (sbentleyaz):

isn't it multiplication

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle (10x)(e^{-x})=\frac{10x}{e^x} }\)

OpenStudy (solomonzelman):

Just algebra, friend .... :)

OpenStudy (sbentleyaz):

oh because the its a negative exponent?

OpenStudy (solomonzelman):

Yes:)

OpenStudy (solomonzelman):

(I would advise, though, that later, when you get some free time read about L'Hospital's rule, and the indeterminate forms.)

OpenStudy (solomonzelman):

Can you set up the limit for (b)?

OpenStudy (sbentleyaz):

yes... so lim as x-> infinity is x^2/e^4x ?

OpenStudy (solomonzelman):

Yes.

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle \lim_{x\to\infty}\frac{x^2}{e^{4x}} }\)

OpenStudy (solomonzelman):

Can you solve this limit?

OpenStudy (sbentleyaz):

OH i think i just figured it out

OpenStudy (solomonzelman):

(If you really want to avoid L'Hospital's rule, you can either do the numerical approach - plug in bigger and bigger values, or you can do the similar thing to what I did before) ...

OpenStudy (solomonzelman):

Figured out, ok ... go for it, of course:)

OpenStudy (sbentleyaz):

i plugged them all into a graphing calculator after i set up the limit.. and i got C (x) =(Lnx)3; g(x) = 1/x.... because this is the only one that does not approach 0

OpenStudy (sbentleyaz):

as it approaches infinity

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle \lim_{x\to \infty}\frac{(\ln x)^3}{x} }\) Like this?

OpenStudy (sbentleyaz):

yes thats exactly what i did

OpenStudy (sbentleyaz):

and then i graphed it

OpenStudy (solomonzelman):

\(\small\color{black}{\displaystyle \lim_{x\to \infty}\frac{(\ln x)^3}{x}=\lim_{x\to \infty}\frac{\frac{d}{dx}(\ln x)^3}{\frac{d}{dx}x}=\lim_{x\to \infty}\frac{3(\ln x)^2\frac{1}{x}}{1}=\lim_{x\to \infty}\frac{3(\ln x)^2}{x} }\) \(\small \color{black}{\displaystyle=\lim_{x\to \infty}\frac{\frac{d}{dx}3(\ln x)^2}{\frac{d}{dx}x}=\lim_{x\to \infty}\frac{6(\ln x)\frac{1}{x}}{1}=\lim_{x\to \infty}\frac{6\ln x}{x} =\lim_{x\to \infty}\frac{\frac{d}{dx}6\ln x}{\frac{d}{dx}x}=\lim_{x\to \infty}\frac{6\frac{1}{x}}{1}=? }\)

OpenStudy (solomonzelman):

I didn't want to disappoint you, but, as I solved by L'Hospital's rule, this limit is indeed going to be 0.

OpenStudy (solomonzelman):

What is f(x) in (d)?

OpenStudy (solomonzelman):

Also, can you tell me the precise f and g for (a)?

OpenStudy (sbentleyaz):

hm thats strange because out of all the ones i graphed, that was the only one that didn't approach 0

OpenStudy (sbentleyaz):

@3mar

OpenStudy (3mar):

Well, I am here. Sorry for being late! I was not here! Still need help?